OpenStudy (saifoo.khan):

2x^2+2y+10. solve for y.

OpenStudy (anonymous):

2x^2 + 2 y+10 = 0

OpenStudy (saifoo.khan):

no, thts wrong.

OpenStudy (anonymous):

2x^2+10=-2y

OpenStudy (anonymous):

y= -x^2 -10

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

-x^2-5

OpenStudy (anonymous):

subtract 2y from both sides, then divide by -2 you end up with y=-x^2-10

OpenStudy (anonymous):

+10/-2 = -10?

OpenStudy (anonymous):

that's right should be -5..thanks!

OpenStudy (saifoo.khan):

whoz right?

OpenStudy (saifoo.khan):

who wants a medal?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

meee

OpenStudy (anonymous):

to solve 2x^2+2y+10 = 0; => x^2+y+5=0; =>y=-(x^2+5); how about that....